OpenStudy (anonymous):

Given cos(theta) = 1/3 and sin(theta)<0, Find cot(theta). (Unit Circle)Given cos(theta) = 1/3 and sin(theta)<0, Find cot(theta). (Unit Circle)@Mathematics

OpenStudy (anonymous):

\[\cos^{-1} (1/3)=70.53°\] \[sin\theta\]=0.942 \[\cot \theta\] =0.353529

OpenStudy (anonymous):

Ok. Can you explain how you got that answer?

OpenStudy (anonymous):

i'm not sure whetr datz ryt:(